OpenStudy (anonymous):

(1+cscx)/(cosx+cotx)i need to show all my work

OpenStudy (anonymous):

u want me to solve or is it closed

OpenStudy (anonymous):

what the answer is

OpenStudy (anonymous):

rewrite in terms of sin and cos... find CDs in the numerator and denominator... it simplfies pretty easily.

OpenStudy (anonymous):

eg. what is 1 +cscx in terms of sin or cos?

OpenStudy (anonymous):

it just says simplify

OpenStudy (anonymous):

awesome! gl!

hartnn (hartnn):

csc x = 1/ sin x so what is 1+csc x = ?

OpenStudy (anonymous):

must. procrastinate. from. diffeq. \[(1+\csc)/(\cos+\cot) = [(\sin/\sin) + (1/\sin) ] / [(\cos*\sin+\cos)/\sin] = [(\sin+1)/\sin] * [\sin/\cos(\sin+1)]\] sin cancels, (sin+1) cancels. end up with 1/cos

OpenStudy (anonymous):

sorry if thats kind of messy lol

hartnn (hartnn):

@irpirate use `\frac{}{}` for fractions `\frac{1+csc}{cos+cot}` will give u \(\frac{1+csc}{cos+cot}\)

OpenStudy (anonymous):

\[ \frac{1+ \frac{1}{\sin}}{\cos+ \frac{\cos}{\sin}} = \frac{\frac{\sin}{\sin}+\frac{1}{\sin}}{\frac{\cos*\sin+\cos}{\sin}} = \frac{\sin+1}{\sin}*\frac{\sin}{\cos(\sin+1)} = \frac{1}{\cos}\] that was a pain lol

hartnn (hartnn):

and u just missed x everywhere :P

OpenStudy (anonymous):

-.-''''''''''''''' i think she'll manage :P

hartnn (hartnn):

@tammyhawatmeh see the solution step by step, and tell us which step did u not get

OpenStudy (anonymous):

i got it thanks guys

OpenStudy (anonymous):

^_^ welcome